[clas_members] [New Logentry] RC Daily Summary August 12, 2022

yprok at jlab.org yprok at jlab.org
Fri Aug 12 10:50:01 EDT 2022

Logentry Text:
Thursday, August 11: 

Day Shift:  Restricted Access  started at 8am (Maintenance Day). 
                 Target material was changed from ND3 to NH3 (15 mm cell). Other work and a safety walk through have been done as well. 

Swing Shift: Hall went into Beam Permit at 4:30pm. Target polarization saturated at about 21%.  Target group decided to change the target material the next day. 
                    Waited for beam the rest of the shift.

Friday, August 12: 

Owl Shift:  Beam came back around 2am. Tuned beam to the tagger, took harp scans, and prepared to take production data. MCC  informed us that  some halo counters were unacceptably high (Hall C leakage problems?). At 4am Crew Chief informed us that they need to replace a fiber card related to the FSD node (see above). FSD card replacement was successful, and beam was ready to be delivered at 5:30 am. Started a run, but experienced DAQ problems; at the same time target polarization went up as soon as it was exposed to the beam, indicating that the material has been under-dosed. Target group advised not to changed the material sample today but to continue irradiating the material until polarization starts to decline, then anneal it. 

Plan for day shift: Optimize raster parameters for the 15 mm cell; take production data. 

We will have the daily RC meeting today at 2:00 PM 

RC Meeting Coordinates:
Meeting ID: 160 867 4253
Passcode: 786506


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4027941
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/clas_members/attachments/20220812/041b5d81/attachment.html>

More information about the clas_members mailing list